### v2

# Emergent Dynamical Translational Symmetry Breaking as an Order Principle for Localization and Topological Transitions

Abstract
Localization transitions represent a fundamental class of continuous phase transitions, yet they occur without any accompanying symmetry breaking. We resolve this by introducing the concept of dynamical translational symmetry (DTS), which is defined not by the Hamiltonian but by the long-time dynamics of local observables. Its order parameter, the time-averaged local translational contrast (TLTC), quantitatively diagnoses whether evolution restores or breaks translational equivalence. We demonstrate that the TLTC universally captures the Anderson localization transition, the many-body localization transition, and topological phase transitions, revealing that these disparate phenomena are unified by the emergent breaking of DTS. This work establishes a unified dynamical-symmetry framework for phases transitions beyond the equilibrium paradigm.
Introduction.— Understanding the nature of phase transitions has long been a central theme in condensed matter physics. Traditionally, the Landau paradigm characterizes continuous phase transitions in terms of spontaneous symmetry breaking and the emergence of local order parameters Landau1 ; Landau2 . However, several important classes of continuous transitions lie beyond this conventional framework, as they do not involve any symmetry breaking. Among the most prominent examples are topological phase transitions, where distinct phases are distinguished not by symmetry but by topological invariants and boundary modes TP1 ; TP2 ; TP3 ; TP4 . The deep understanding of such transitions developed over the past few decades has profoundly expanded our view of quantum matter. In parallel, localization transitions pose an equally profound challenge to the Landau framework. These include (i) the Anderson localization transition, where single-particle eigenstates evolve continuously from extended to localized as disorder increases AL1 ; AL2 ; AL3 ; AL4 , and (ii) the many-body localization (MBL) transition, where interacting disordered systems exhibit a breakdown of ergodicity and thermalization MBL1 ; MBL2 ; MBL3 ; MBL4 ; MBL5 . Both represent continuous phase transitions that occur without any accompanying symmetry breaking.
The fundamental nature of localization transitions is directly reflected in profound alterations of the system’s dynamical behavior. In the extended or thermal regime, even when the Hamiltonian contains disorder or quasiperiodic potentials, the system exhibits ballistic diffusion identical to a periodic system: an initially localized excitation spreads throughout the lattice, dynamically restoring spatial homogeneity. In contrast, in the localized regime, the system retains memory of its initial configuration, resulting in persistent spatial confinement. This observation motivates a unifying perspective that focuses not on the static symmetries of the Hamiltonian, but on the symmetries emerging from the system’s long-time dynamical evolution. Consequently, prolonged dynamical evolution may give rise to emergent translational symmetry absent in the original Hamiltonian, a feature we term dynamical translational symmetry (DTS). Unlike conventional order parameters associated with static symmetry breaking, DTS captures the translational behavior emerging in the dynamics of local observables. Its breakdown signals the failure of dynamical homogenization, thereby distinguishing localized and extended regimes. Remarkably, we further find that DTS breaking can also characterize topological phase transitions.
In this work, we establish DTS breaking as a dynamical order principle for localization and topological transitions. Preservation of DTS corresponds to ergodic or extended dynamics, while its breaking signals localization, memory retention, or boundary confinement. Through the quantitative framework of the time-averaged local translational contrast (TLTC), we demonstrate that Anderson localization, MBL, and topological phase transitions can all be consistently interpreted as distinct manifestations of the same underlying dynamical symmetry principle, thereby extending the Landau paradigm of symmetry breaking to localization and topological physics.
Dynamical Translational Symmetry.— For a static lattice Hamiltonian $H$, ordinary translational symmetry requires $T_{a},H]=0$, where $T_{a}$ is the lattice translation operator that shifts all sites by $a$ lattice spacings. This commutation ensures that the time-evolution operator $U(t)=e^{-iHt}$ also satisfies $U(t),\mathcal{T}_{a}]=0$, so the dynamics remain translationally invariant at all times. In the presence of disorder or quasiperiodicity, this static symmetry is explicitly broken at the Hamiltonian level. Nevertheless, translational invariance may emerge dynamically: after long-time evolution, local observables can become effectively position-independent, even though $H$ itself lacks translation invariance. We refer to this emergent invariance as DTS, which characterizes whether the long-time evolution restores translational equivalence across lattice sites. Operationally, DTS implies that two identical local probes placed at different sites record identical time-averaged signals, while any deviation between them signifies persistent spatial inhomogeneity and the breaking of DTS.
To quantify the degree of DTS breaking, we introduce the TLTC, defined as
$\mathcal{C}_{a}^{(O)}(T_{f},T_{i},j)=\frac{1}{T_{f}-T_{i}}\int_{T_{i}}^{T_{f}}\Big|\big\langle\psi(0)\big|\Big(U^{\dagger}(t)\,O_{j}\,U(t)-\mathcal{T}_{a}[U^{\dagger}(t)\,O_{j}\,U(t)]\Big)\big|\psi(0)\big\rangle\Big|^{2}dt,$ (1)
where $O_{j}$ is a local observable at site $j$, $|\psi(0)\rangle$ is a normalized initial state (typically a localized wave packet), and $\mathcal{T}_{a}[X]\equiv T_{a}^{\dagger}XT_{a}$ The parameters $T_{i}$ and $T_{f}$ denote the lower and upper limits of the time window used for averaging.
The TLTC measures the time-averaged deviation between the local expectation value of an observable and that of its translated counterpart during the dynamical evolution. For an arbitrary initial state $|\psi_{0}\rangle$, the long-time average of a local observable $O_{j}$ is defined as
$\overline{O_{j}}=\frac{1}{T}\int_{0}^{T}\langle\psi(0)|U^{\dagger}(t)O_{j}U(t)|\psi(0)\rangle\,dt,$
If the TLTC in Eq. (1) approaches zero in the long-time limit, $\lim_{T_{f}\to\infty}\mathcal{C}_{a}^{(O)}(T_{f},T_{i},j)=0$, then $\overline{O_{j}}=\overline{O_{j+a}}$, implying that long-time averaged local observables become translationally equivalent, meaning that the system exhibits emergent spatial homogeneity. Conversely, if $\mathcal{C}_{a}^{(O)}$ saturates to a finite value, $\overline{O_{j}}\neq\overline{O_{j+a}}$, it signals persistent dynamical inhomogeneity and hence broken DTS. In the End Matter, we rigorously prove that in ergodic (thermalizing) phases, the TLTC always vanishes in the long-time limit, i.e., $\lim_{T_{f}\to\infty}\mathcal{C}_{a}^{(O)}=0$ for any local observable $O_{j}$. This indicates that, although the microscopic Hamiltonian $H$ may explicitly break translational symmetry due to disorder or quasiperiodicity, translational invariance dynamically emerges during long-time evolution in ergodic systems.
The TLTC thus functions as a dynamical order parameter, providing a quantitative criterion for the emergence or breaking of DTS. It vanishes in the symmetric (extended or ergodic) phase and acquires a finite value in the symmetry-broken (localized or dynamically confined) phase. This behavior directly parallels the role of static order parameters in Landau’s paradigm. Importantly, the TLTC is basis-independent and universally applicable, ranging from single-particle to interacting localization and topological systems. As we shall demonstrate below, the breaking of DTS provides a unifying language for describing Anderson localization, MBL, and topological transitions within a single dynamical framework.
In Eq. (1), the choice of the averaging window $(T_{i},T_{f})$ is flexible; typically one requires $T_{f}\gg T_{i}$ to ensure convergence. For finite systems, it is not necessary to take the strict limit $T_{f}\to\infty$; a sufficiently long but finite evolution time already captures the asymptotic behavior. In practice, for both analytical convenience and numerical simulations, we fix $T_{i}=0$ and denote the upper limit simply by $T_{f}=T$. Furthermore, it is unnecessary to evaluate Eq. (1) for every lattice site $j$. Since DTS concerns translational equivalence, it suffices to monitor a representative site, typically the one where the local excitation or contrast is maximal at $t=0$. For example, in the study of Anderson localization, we consider an initial wave packet $|\psi(0)\rangle$ localized at site $i_{0}$, and evaluate the corresponding $\mathcal{C}_{a}^{(O)}(T,0,i_{0})$. If this quantity tends to zero in the long-time limit, $\mathcal{C}_{a}^{(O)}(T,0,j)\to 0$ for all $j$, signifying complete restoration of DTS. Without loss of generality, we set $a=1$ in the following and denote $\mathcal{C}_{a}^{(O)}(T_{f},T_{i},j)$ simply as $\mathcal{C}_{a}^{(O)}(T)$.
Anderson localization transition.— We first illustrate that DTS breaking can characterize the Anderson localization transition using the Aubry-André (AA) model AA , a paradigmatic realization of localization transition in quasiperiodic systems. The single-particle Hamiltonian is
$H_{\text{AA}}=-J\sum_{i=1}^{L-1}(c_{i}^{\dagger}c_{i+1}+\mathrm{h.c.})+\lambda\sum_{i=1}^{L}\cos(2\pi\beta i+\phi)\,n_{i},$ (2)
where $c_{i}^{\dagger}$ ($c_{i}$) creates (annihilates) a particle on site $i$, $J$ is the hopping amplitude, $\lambda$ controls the strength of the quasiperiodic potential, $\beta$ is an irrational number, and $\phi$ is the initial phase. This model exhibits a self-dual localization transition at $\lambda_{c}/J=2$: for $\lambda<\lambda_{c}$, all eigenstates are extended, while for $\lambda>\lambda_{c}$ they become exponentially localized. Without loss of generality, in the following calculations we fix $\beta=(\sqrt{5}-1)/2$ and $\phi=0$.
To probe DTS, we compute the TLTC,
$\mathcal{C}_{a}^{(P)}(T)=\frac{1}{T}\int_{0}^{T}\big|P_{i_{0}}(t)-P_{i_{0}+a}(t)\big|^{2}\,dt,$ (3)
where $P_{i}(t)=|\psi_{i}(t)|^{2}$ is the instantaneous local probability, and $i_{0}$ denotes the initial position of the wave packet. A vanishing $\mathcal{C}_{a}^{(P)}$ indicates that the long-time wave function becomes translationally uniform, signaling preserved DTS, whereas a finite value implies its breaking due to localization.
Figure 1(a) shows the time evolution of $\mathcal{C}_{a}^{(P)}(T)$ for representative potential strengths $\lambda/J=1,2,3$. In the extended regime ($\lambda/J=1$), $\mathcal{C}_{a}^{(P)}(T)$ decays algebraically toward zero, indicating dynamical restoration of translational homogeneity. At the critical point ($\lambda/J=2$), the decay slows down and becomes nonmonotonic, reflecting critical fluctuations in the spreading process. In the localized regime ($\lambda/J=3$), $\mathcal{C}_{a}^{(P)}(T)$ rapidly saturates to a finite value, demonstrating the breaking of DTS. The inset shows the long-time averaged local probability distribution $\overline{P_{i}}=\frac{1}{T}\int_{0}^{T}P_{i}(t)\,dt$, which becomes uniform in the extended phase while remaining exponentially localized in the localized phase. We fix $T=1000$ and plot the long-time value $\mathcal{C}_{a}^{(P)}(T=1000)$ as a function of $\lambda/J$ in Fig. 1(b), which clearly reveals that $\mathcal{C}_{a}^{(P)}(T=1000)$ changes from zero to a finite value as the system transitions from the extended to the localized regime. The TLTC thus serves as a dynamical order parameter for the Anderson localization transition. Unlike static indicators such as the inverse participation ratio, which rely on eigenstate properties, the TLTC directly employs local observables to capture the restoration or breakdown of translational symmetry during long-time evolution, providing a unified dynamical-symmetry perspective on the localization transition.
Figure 1: (a) Time evolution of the TLTC indicator $\mathcal{C}_{a}^{(P)}(T)$ for representative $\lambda/J=1,2,3$. The inset shows the corresponding long-time averaged site probabilities $\overline{P_{i}}$. (b) Long-time value $\mathcal{C}_{a}^{(P)}(T=1000)$ versus $\lambda/J$. Here we fix $J=1$, $L=610$, $i_{0}=L/2$, $a=1$, use a time integration step of $dt=0.2$, and employ open boundary conditions (OBC).
Figure 2: (a) Time evolution of the TLTC, $\mathcal{C}_{a}^{(P)}(T)$ (solid), and its site-averaged counterpart $\overline{\mathcal{C}}_{a}^{(P)}(T)$ (dashed) in the interacting AA model. For $\lambda/J=1$ (red), both quantities decay to zero, indicating ergodic dynamics with restored DTS, whereas for $\lambda/J=4$ (blue), they saturate to finite values, reflecting MBL behavior. (b) Comparison between the instantaneous and time-averaged TLTC, $\mathcal{C}_{a}^{(P)}(t)$ and $\mathcal{C}_{a}^{(P)}(T)$, and the density imbalance, $\mathcal{I}(t)$ and $\mathcal{I}(T)$, as functions of the quasiperiodic potential strength $\lambda/J$. All four quantities change from vanishing to finite values at approximately the same critical $\lambda/J$. Here we fix $J=1$, $V=1$, $L=14$, $N=7$, $i_{0}=7$, $a=1$, $dt=0.5$, and use OBC.
Many-body localization transition.— The framework of DTS breaking extends naturally to interacting systems. We next investigate DTS breaking in the context of MBL, where ergodicity and thermalization break down. In contrast to the single-particle AA model, where localization arises from quasiperiodic potential modulation, the MBL transition emerges from the interplay between disorder or quasiperiodicity and interparticle interactions. To study this, we consider the interacting AA model,
$H_{\text{AA+int}}=H_{\text{AA}}+V\sum_{i=1}^{L-1}n_{i}n_{i+1},$ (4)
where $V$ is the nearest-neighbor interaction strength. The system undergoes an ergodic-to-MBL transition as $\lambda/J$ increases for a fixed $V$ AAinter .
We fix the system size to $L=14$ and the particle number to $N=7$. The initial state is chosen as a charge-density-wave configuration, where particles occupy odd lattice sites. We employ Eq. (3) to characterize the distinct dynamical features of different many-body phases, with $P_{i}(t)=\langle\Psi(t)|n_{i}|\Psi(t)\rangle$, where $|\Psi(t)\rangle$ denotes the many-body wavefunction at time $t$. Again taking $a=1$, the reference site $i_{0}$ can be chosen arbitrarily, since all sites satisfy the condition of maximal initial contrast. For comparison, we also introduce a site-averaged TLTC:
$\overline{\mathcal{C}}_{a}^{(P)}(T)=\frac{1}{TL}\int_{0}^{T}\!\!dt\,\sum_{i=1}^{L}\big|P_{i}(t)-P_{(i+a)\,\mathrm{mod}\,L}(t)\big|^{2},$ (5)
Figure 2(a) shows the evolution of both $\mathcal{C}_{a}^{(P)}(T)$ and $\overline{\mathcal{C}}_{a}^{(P)}(T)$ over time. For $\lambda/J=1$, both quantities decay to zero, indicating that local densities become homogeneous, corresponding to the ergodic (thermalizing) phase. For $\lambda/J=4$, they decay slowly and saturate to finite nonzero values, reflecting the onset of MBL and the associated breaking of DTS. Throughout the evolution, $\mathcal{C}_{a}^{(P)}(T)$ and $\overline{\mathcal{C}}_{a}^{(P)}(T)$ remain nearly identical, demonstrating that the localization properties of an interacting system can also be faithfully inferred from the dynamics of only two lattice sites.
To benchmark against the standard experimental diagnostic of many-body localization, we compare the instantaneous and time-averaged TLTC, $\mathcal{C}_{a}^{(P)}(t)=|P_{i_{0}}(t)-P_{i_{0}+a}(t)|^{2}$ and $\mathcal{C}_{a}^{(P)}(T)$ (defined in Eq. (3)), with the experimentally measurable density imbalance $\mathcal{I}(t)$ and its time average $\mathcal{I}(T)$, defined as MBL4
$\mathcal{I}(t)=\frac{1}{N}\sum_{i=1}^{L}(-1)^{i}\,\langle n_{i}(t)\rangle,\quad\mathcal{I}(T)=\frac{1}{T}\int_{0}^{T}\mathcal{I}(t)\,dt.$ (6)
A vanishing imbalance, $\mathcal{I}(t)\!\to\!0$, indicates ergodic dynamics, whereas a finite saturation value signals MBL. Figure 2(b) shows the variation of these quantities as a function of the quasiperiodic potential strength $\lambda$. One observes that $\mathcal{C}_{a}^{(P)}(t)$, $\mathcal{C}_{a}^{(P)}(T)$, $\mathcal{I}(t)$, and $\mathcal{I}(T)$ all transition from vanishing to finite values at approximately the same critical $\lambda/J$, demonstrating that the TLTC faithfully captures the ergodic-to-MBL transition. Unlike the imbalance, however, measuring $\mathcal{C}_{a}^{(P)}$ requires monitoring only two lattice sites with initially contrasting occupations, such as one occupied and one empty site, regardless of the detailed form of the prepared initial state. This simplicity makes TLTC-based characterization particularly favorable for experimental studies of MBL transitions. The TLTC thus acts as a unified dynamical order parameter for both Anderson and MBL transitions, quantitatively linking ergodic spreading to localized memory retention.
Topological transition.— The concept of DTS breaking can be further extended to describe topological phase transitions, where the presence of boundary modes intrinsically breaks DTS. We illustrate this connection using the Su-Schrieffer-Heeger (SSH) model SSH , which is given by
$H_{\text{SSH}}=-\sum_{i=1}^{L-1}\big[J_{1}\,c_{2i-1}^{\dagger}c_{2i}+J_{2}\,c_{2i}^{\dagger}c_{2i+1}+\mathrm{h.c.}\big],$ (7)
where $J_{1}$ and $J_{2}$ denote alternating hopping amplitudes. The system is topologically nontrivial for $J_{2}/J_{1}>1$ and trivial for $J_{2}/J_{1}<1$. In the topological regime, open boundaries support exponentially localized zero-energy edge states.
We initialize a single-particle wave packet localized at the boundary site, $\psi_{i}(0)=\delta_{i,1}$, and monitor its time evolution. We again evaluate the TLTC defined in Eq. (3), taking $i_{0}=1$ and $a=1$. As shown in Fig. 3(a), in the trivial phase ($J_{2}/J_{1}=0.5$), the TLTC $\mathcal{C}_{a}^{(P)}(T)$ rapidly decays to zero, indicating dynamical translational equivalence across the lattice. This behavior reflects that the initially localized excitation spreads throughout the system, dynamically restoring spatial homogeneity and thereby preserving DTS. In contrast, in the topological phase ($J_{2}/J_{1}=1.5$), $\mathcal{C}_{a}^{(P)}(T)$ saturates to a finite nonzero value, signaling boundary-induced DTS breaking. The edge-localized mode remains confined near the boundary due to topological protection, resulting in persistent spatial asymmetry even after long-time evolution. We fix $T=1000$, and Fig. 3(b) shows $\mathcal{C}_{a}^{(P)}(T=1000)$ as a function of the hopping ratio $J_{2}/J_{1}$. As the system evolves from the topologically trivial to the nontrivial regime, $\mathcal{C}_{a}^{(P)}(T=1000)$ increases from zero to a finite value. This demonstrates that the TLTC serves as an effective order parameter for the topological transition, reflecting the breaking of DTS localized at the boundary.
Figure 3: (a) Time evolution of the TLTC $\mathcal{C}_{a}^{(P)}(T)$ in the SSH model for $J_{2}/J_{1}=0.5$ (red) and $J_{2}/J_{1}=1.5$ (blue), with the particle initially localized at the boundary site. (b) The long-time value $\mathcal{C}_{a}^{(P)}(T{=}1000)$ as a function of $J_{2}/J_{1}$. Here we fix $J_{1}=1$, $L=600$, $i_{0}=1$, $a=1$, $dt=0.5$, and use OBC.
Conclusion and discussion.— We have introduced the concept of dynamical translational symmetry (DTS) and formulated the time-averaged local translational contrast (TLTC) as its quantitative measure. The preservation of DTS corresponds to ergodic or extended dynamics, while its breaking signifies localization, memory retention, or boundary confinement. Although the microscopic mechanisms differ between localized states in disordered systems and topological edge states in topological phases, their behaviors can be consistently interpreted as the emergence of DTS breaking arising from nonergodic dynamical evolution. The TLTC therefore serves as a dynamical order parameter, analogous to static order parameters in Landau theory, but defined through the long-time evolution of local observables.
The concept of DTS can be naturally extended to driven and open quantum systems explainopen , where emergent dynamical symmetry may interplay with Floquet synchronization or dissipation-induced ordering. Moreover, because the TLTC relies solely on local observables and time averaging, it can be directly measured in various experimental platforms, including ultracold atoms, superconducting quantum circuits, and photonic simulators. We anticipate that DTS and its breaking will provide a versatile framework for characterizing nonequilibrium quantum phases and for understanding the emergence of dynamical order in complex quantum systems.
Acknowledgements.
This work is supported by National Key R&D Program of China under Grant No.2022YFA1405800, the Key-Area Research and Development Program of Guangdong Province (Grant No.2018B030326001), Guangdong Provincial Key Laboratory(Grant No.2019B121203002).